Problem #4 VCE = 8 +0.7=6 8.7V Rsig RL=0 VA 150 + 10 (1+12 170 V ſo=100 15V o www 0.1 uF NCE Vi No & skr ਫੀਸਦ · e) End Avo Risolve
Answer #1

Given p=100 Given Vee 38.7 V VA = 1700 Nce = 8.7V Here Vi= Vcc VCE 15= Vcc R, & VEL ²3k2 Vo = Vo=15-8.7 No=67 volts: Vo =IE (General Design Rules, Generally tu bias in stable LORth < (3+1). RE R,R ~ (loot! 13 km) RAR 10x Rif = 30.3 RIR = 30-3 (R+R2)small sig © signal parameters 2:08 MA = 0.08 Alv Y 26 mv fm = I Xx = ß = 1.25 KM 0.0% gm Vo= VA 170 =81.73 km 2:08MA IL d) IEes Rout= Il a So IIRE 3+1 125 81.73113k.ll 1004) => Rout = 12-321 ki z Ya+(+1)(ell ro) 1·25+(* [ 3 || 8173) Ri= 243.526 (Bt{